An athlete completes one round of a circular track of diameter 200 m in 40 s. What will be the distance covered and the displacement at the end of 2 minutes 20 s?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Diameter of circular track (D) = 200 m
Radius of circular track (r) = 200 / 2=100 m
Time taken by the athlete for one round (t) = 40 s
Distance covered by athlete in one round (s) = 2π r
= 2 * ( 22 / 7 ) * 100
Speed of the athlete (v) = Distance / Time
= (2 x 2200) / (7 x 40)
= 4400 / 7 × 40
So,Distance covered in 140 s = Speed (s) × Time(t)
= 4400 / (7 x 40) x (2 x 60 + 20)
= 4400 / ( 7 x 40) x 140
= 4400 x 140 /7 x 40
= 2200 m
Number of rounds in 40 s =1 round
Number of rounds in 140 s =140/40
=3 ½
Therefore,Displacement of the athletewith respect to initial position at x= xy
= Diameter of circular track
= 200 m

Thanks